OpenStudy (anonymous):

solve by factoring x^2+x=12

OpenStudy (anonymous):

x=-4,3

OpenStudy (anonymous):

(x +4)(x - 3) = 0 x = 3 , -4

OpenStudy (anonymous):

\[x^2+x-12=0\] \[(x+4)(x-3)=0\] \[x=-4 or x=3\]